OpenStudy (angela210793):

It's given: y=-5+7cos(3a-9).From which of the following values the function takes maximum: a) -177 degrees b) -123 degrees c) 0 degrees d) 111 degrees

OpenStudy (amistre64):

0 wins the plug and play

OpenStudy (angela210793):

O.o I thought it should be cos(3a-9)=1 and u had to find a but I didn't get 0.... how did u do tht?Explain pls????????

OpenStudy (amistre64):

i just plugged in the values of the answers into the equaton to see which one was the biggest

OpenStudy (amistre64):

another way is to just go by the 3a-9 part and see which angles higher on the unit circle

OpenStudy (angela210793):

wouldn't it b 3*0-9=-9? how do u find it's cos? O.o

OpenStudy (amistre64):

the cos of any angle gets smaller as you move from 0 up to 90; and also from 0 to -90; so that largest 'value in that will get your biggest number right?

OpenStudy (amistre64):

cos(-9) is bigger than cos(10) by reason that 10 produces a steeper angle that has a cosine value closer to 0 than to 1

OpenStudy (angela210793):

oh......I got it....Thanks :):):) smth else pls....what does that guru mean?:)))

OpenStudy (amistre64):

3(-177)-9 -531-9 = -540; spin out 360 to get it to; -180 cos(-180) = -1 3(-123)-9 -369-9 = -378; spin out a 360 to get it to; -18 cos(-18) is between 0 and 1 3(0)-9 = -9 cos(-9) is closer to 1 than cos(-18) 3(111)-9 333-9 = 324; to keep things the same spin out 360 to get; -36 cos(-36) is even farther away from 1 than cos(-9) so the largest value for the cos(...) part is obtained at a=0 which then gives you the largest value for the entire equation

OpenStudy (angela210793):

I've always hated trigonometry :(:(:( Thank you veeeeeeryyyy much :):)

OpenStudy (amistre64):

this might splain it better :)

OpenStudy (angela210793):

Yes it does :):) ^_^ Thnx :):)